What is the value of w in the equation 3/4w + 8= 1/3w-7?
Answer by rfer(16322)

3/4w-1/3w=-7-8
9/12w-4/12w=-15
5/12w=-15
w=-15/(5/12)
w=-15*12/5
w=-36